OpenStudy (anonymous):

the first outcome is one of X, Y, and Z, and the second outcome in each case is one of A and B. Pr[X] = [ 4/7] Pr[Y] = [ 1/7] Pr[Z] = [ 2/7] Pr[A | X] = [ 2/5] Pr[B | X] = [ 3/5] Pr[A | Y] = [ 1/3] Pr[B | Y] = [ 2/3] Pr[A | Z] = [ 5/8] Pr[B | Z] = [ 3/8] Find the missing probability: Pr[(X or Y) | B]

OpenStudy (kirbykirby):

First, recognize from the rule for unions that: \(P(X~or~Y|B)=P(X \cup Y|B)=P(X|B)+P(Y|B)-P(X \cap Y|B)\). But, since the first event can only X or Y or Z , you can't have two of these events occurring simultaneously meaning that \(P(X \cap Y|B)=0\) Since X, Y and Z form a partition over the sample space, you can use Bayes' Theorem, and find that \[P(X|B)=\frac{P(B|X)P(X)}{P(B|X)P(X)+P(B|Y)P(Y)+P(B|Z)P(Z)} \\ = \frac{(3/5)(4/7)}{(3/5)(4/7)+(2/3)(1/7)+(3/8)(2/7)}\] Similarly, \[P(Y|B)=\frac{P(B|Y)P(Y)}{P(B|X)P(X)+P(B|Y)P(Y)+P(B|Z)P(Z)} \\ = \frac{(2/3)(1/7)}{(3/5)(4/7)+(2/3)(1/7)+(3/8)(2/7)} \]

OpenStudy (kirbykirby):

then just add those 2 probabilities together

OpenStudy (anonymous):

This was correct, great explanation - thank you so much!

OpenStudy (kirbykirby):

awesome :)